1.Diego is considering eliminating the West region because an internally generated report suggests the region’s total gross margin in the first year of operations was $50,000 less than its traceable fixed selling and administrative expenses. Diego believes that if it drops the West region, the East region's sales will grow by 5% in Year 2. Using the contribution approach for analyzing segment profitability and assuming all else remains constant in Year 2, what would be the profit impact of dropping the West region in Year 2?

2. Assume the West region invests $38,000 in a new advertising campaign in Year 2 that increases its unit sales by 20%. If all else remains constant, what would be the profit impact of pursuing the advertising campaign?Required information The Foundational 15 (LO6-1, LO6-2, LO6-3. LO6-4. LO6-51 [The following information applies to the questi
Answer #1
Variable cost per unit 47 =27+12+3+5
Unit Contribution margin 30 =77-47
1
Loss in Contribution margin of West -300000 =10000*30
Avoidable fixed selling and administrative expenses 220000
Increase in Contribution margin of East 49500 =33000*5%*30
Net change in profits -30500
Profits will decrease by 30500
2
Increase in Contribution margin of West 60000 =10000*20%*30
Less: Advertising expense -38000
Net change in profits 22000
Profits will increase by 22000